60 ΚΩ 30 ΕΩ 50 ΚΩ 20 ΚΩ v0 / vi = ?
Answer #1

The circuit has two opamps. It is marked as 1 and 2 in the figure. The output of teh first op- amp is marked as V_{o1} and the output of the second op - amp is marked as V_{o2} as shown in figure below

60 ΚΩ 30 ΚΩ Vo1 1 50 ΚΩ RL 20 ΚΩ ν 2 V2

The op amp on the top that is marked 1 is configured as a non - inverting amplifier. The output of this amplifier will be

V_{o1}=\Big( 1 + \frac{60\times 10^3}{30\times 10^3} \Big)\times v_i

V_{o1}=3v_i

The op amp on the bottom that is marked 2 is configured as an inverting amplifier. The output of this amplifier will be

V_{o2}=-\Big( \frac{50\times 10^3}{20\times 10^3} \Big)\times v_i

V_{o2}=-2.5v_i

The output across the load resistor R_L is the difference between these two outputs. So

v_o = V_{o1}-V_{o2}

v_o = 3v_i- -2.5v_i

v_o = 5.5v_i

So the output input ratio is

\frac{v_o}{v_i}= 5.5